Reduce the equation Zxx - (1+y)2 Zyy = 0 to canonical form


1
Expert's answer
2021-06-07T19:13:13-0400

from the given equation,

R=1, S=0,T=-(1+y)2

By using S2-4RT=0-4(-(1+y)2 =4(1+y)2 >0.

Therefore, the given equation is a hyperbolic partial Differential equation.

And its canonical form is given by

"\frac{\partial^2z}{\partial u\partial v}=0"